OpenStudy (anonymous):

find the equation in point-slope form of the line that is the perpendicular bisector of the segment between (16,-4) and (-2,-76)

OpenStudy (aayushi.somani):

point slope from is- (y-y1)=m(x-x1) soo.. y+4=m(x-16) but for this u need m i.e. slope.. \[m=\frac{ y2-y1 }{ x2-x1 }\] soo it comes out to be 4.4 now u just need to do is place the values in the above equation and u will be able to find the equation..